### v1

# Unbiased Learning to Rank Meets Reality: Lessons from Baidu’s Large-Scale Search Dataset

Abstract.
Unbiased learning-to-rank (ULTR) is a well-established framework for learning from user clicks, which are often biased by the ranker collecting the data. While theoretically justified and extensively tested in simulation, ULTR techniques lack empirical validation, especially on modern search engines. The dataset released for the WSDM Cup 2023, collected from Baidu’s search engine, offers a rare opportunity to assess the real-world performance of prominent ULTR techniques. Despite multiple submissions during the WSDM Cup 2023 and the subsequent NTCIR ULTRE-2 task, it remains unclear whether the observed improvements stem from applying ULTR or other learning techniques. We revisit and extend the available experiments. We find that unbiased learning-to-rank techniques do not bring clear performance improvements, especially compared to the stark differences brought by the choice of ranking loss and query-document features. Our experiments reveal that ULTR robustly improves click prediction. However, these gains in click prediction do not translate to enhanced ranking performance on expert relevance annotations, implying that conclusions strongly depend on how success is measured in this benchmark.
submissionid: 4266journalyear: 2024copyright: rightsretainedconference: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 14–18, 2024; Washington D.C., USAbooktitle: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’24), July 14–18, 2024, Washington D.C., USA
1. Introduction
Historically, the field of learning-to-rank employed human experts to annotate the relevance of search results to create training data for ranking models (Qin and Liu, 2013; Dato et al., 2016, 2022). As expert labeling is expensive (Chapelle and Chang, 2011), infeasible in certain applications (Wang et al., 2016), and potentially misaligned with user preferences (Sanderson et al., 2010), many practitioners seek to leverage implicit user feedback, often collected in the form of clicks. However, clicks are usually a biased signal of relevance, as they often depend on the position of an item (Joachims et al., 2005), its surrounding items (Craswell et al., 2008; Zhuang et al., 2021), or even the user’s trust in the system to place relevant items on top (Joachims et al., 2005; Agarwal et al., 2019a; Vardasbi et al., 2020b). Over the years, the field of unbiased learning-to-rank (ULTR) has proposed methods to mitigate biases when training ranking models on click data.
ULTR methods, especially in academic research, are often evaluated in synthetic or semi-synthetic setups (Joachims et al., 2017; Ai et al., 2018a; Vardasbi et al., 2020b, a; Oosterhuis and de Rijke, 2021). The most common semi-synthetic setup goes back to the seminal work by Joachims et al. (2017). The setup uses learning-to-rank datasets from web search engines containing features and expert judgments of each query-document pair (Chapelle and Chang, 2011; Qin and Liu, 2013; Dato et al., 2016) and simulates clicks and biases based on synthetic user models (e.g., the position-based model (PBM(Richardson et al., 2007; Craswell et al., 2008) or the cascade model (Craswell et al., 2008)).
Figure 1. Original comparison on Baidu-ULTR, showing four ULTR methods not significantly outperforming a naive method not correcting for position bias. Data from (Zou et al., 2022, Table 6) and visualization by us.
While the semi-synthetic simulation setup has demonstrated the effectiveness of many ULTR methods, it is often unclear to academic researchers how these methods fare in a real-world setup. Recently, Zou et al. (2022) released Baidu-ULTR, an extensive real-world dataset for web search. It comprises over $1.2$ billion user sessions with click data and $397,572$ annotated query-document pairs for evaluation. The dataset contains rich user feedback, including clicks, dwell time, whether a document was scrolled off-screen, and whether the user returned to the result page after clicking an item. In contrast to the classic LTR datasets used in semi-synthetic simulation setups, the dataset does not contain pre-computed ranking features, but the original search query, title, and abstract of each result (for privacy reasons in tokenized form, with a private vocabulary). Therefore, the dataset enables training and evaluating transformer-based rankers, such as MonoBERT (Nogueira et al., 2019) or MonoT5 (Pradeep et al., 2021).
Zou et al. (2022) train a MonoBERT ranker and provide baseline results of this model fine-tuned on four standard ULTR methods: inverse propensity scoring (IPS(Wang et al., 2016; Joachims et al., 2017), regression expectation-maximization (RegressionEM(Wang et al., 2018), the dual learning algorithm (DLA(Ai et al., 2018a), and pairwise debiasing (PairD(Hu et al., 2019). Interestingly, Zou et al. (2022) find that these prominent ULTR methods do not perform significantly better than a naive BERT model trained on clicks without any bias correction, [^1] Zou et al. (2022) do find ULTR to be beneficial on frequent head queries compared to infrequent tail queries. as we report in Figure 1. In subsequent work, the Baidu-ULTR dataset was used at multiple ranking competitions (the WSDM Cup 2023 (Zou et al., 2023b, a) and NTCIR’s ULTRE-2 (Niu et al., 2023)), in which multiple teams reported noticeably higher ranking performance, with and without ULTR techniques. It therefore remains unclear whether the observed improvements were due to unbiased learning-to-rank, and ultimately whether ULTR techniques improve performance on this dataset.
In this work, we reproduce and extend the experiments by Zou et al. (2022), in light of several considerations:
  1. (1)
    The initial findings of ULTR methods bringing barely any benefits on the largest real-world dataset available for ULTR have substantial implications for the field and warrant more scrutiny.
  2. (2)
    Three competitions have been conducted on the Baidu-ULTR dataset, with participants reporting vastly improved ranking results (although not necessarily using ULTR). A preliminary experiment by us revealed that a random ranking of the annotated test dataset achieves a DCG@10 $\approx 6.69$, substantially higher than all baselines in Figure 1 and thus, hinting at potential problems in the original experiments.
  3. (3)
    According to the official codebase that comes with the Baidu-ULTR dataset, [^2] https://github.com/ChuXiaokai/baidu_ultr_dataset/ the original comparison focuses on pointwise versions of the compared ranking methods (except for pairwise debiasing). Therefore, it remains open how pointwise, pairwise, and listwise ULTR methods compare on this dataset.
  4. (4)
    And lastly, a closer look at the Baidu-ULTR dataset and the original codebase revealed some disputable design decisions. IPS, for example, requires an estimate of position bias on the given dataset, while the creators of the Baidu-ULTR dataset confirmed using a hardcoded position bias from the ULTRA library in their experiments (Tran et al., 2021). We also found that placeholder documents make up for almost $20\%$ of the Baidu-ULTR dataset (see Section 3), with an unknown impact on the trained models.
By revisiting the experiments reported by Zou et al. (2022) and the cups’ participants, we address the following research questions:
  1. (RQ1)
    Does unbiased learning-to-rank improve performance on the Baidu-ULTR dataset over naive, non-debiasing models?
  2. (RQ2)
    How do ULTR methods fare against each other, and how do ranking losses and input features affect their performance?
  3. (RQ3)
    Can ULTR methods be applied during language model training, and do they bring improved performance?
We answer (RQ1) negatively as we cannot find robust and significant improvements in a fair comparison between ULTR-based and naive methods. We address (RQ2) by reporting minor but existing differences between ULTR methods and by showcasing large and reliable differences due to ranking losses and input features. Regarding (RQ3), we find that while certain ULTR methods improve over their naive counterpart if applied during language model training, the interactions of ULTR and transformer-based models are not well understood and can even lead to decreased performance.
Besides answering our research questions, our contributions are:
\@afterheading
2. Related Work
2.1. Unbiased Learning to Rank
The field of unbiased learning-to-rank can be broadly divided into click modeling and counterfactual learning-to-rank. Click modeling encodes assumptions on user behavior into probabilistic models (Craswell et al., 2008; Dupret and Piwowarski, 2008; Chapelle and Zhang, 2009). Assumed effects, such as position bias or item relevance, are represented as (latent) variables that are jointly inferred via maximum likelihood estimation (Chuklin et al., 2015). Notable examples include the position-based model (PBM(Richardson et al., 2007; Craswell et al., 2008), which assumes that users click only on positions they examine and items they find relevant, and the cascade model (Craswell et al., 2008), which assumes that users scroll from top to bottom, click on the first relevant item, then leave the page. In recent years, more complex and flexible click models have been proposed using neural architectures, semantic embeddings, and bias features beyond position (Borisov et al., 2016; Chen et al., 2020; Zhuang et al., 2021; Yan et al., 2022). We implement two neural versions of the original PBM: RegressionEM (Wang et al., 2018) and an additive two-tower model popular in industry applications (Guo et al., 2019; Zhang et al., 2023; Zhao et al., 2019).
The counterfactual learning-to-rank community has historically employed simpler user models such as the PBM (Wang et al., 2016; Joachims et al., 2017) or the affine model for trust bias (Vardasbi et al., 2020b) and focused on training more effective ranking models, mitigating biases with inverse propensity scoring (Wang et al., 2016; Joachims et al., 2017; Oosterhuis and de Rijke, 2021). Our work includes an extended binary cross-entropy loss (Bekker et al., 2019; Saito et al., 2020) as a pointwise IPS baseline and an extended softmax loss as a listwise IPS baseline (Bruch et al., 2019; Ai et al., 2018a). Both methods require position bias estimations, for which we release the implementation of three intervention harvesting methods (more in Section 3.2). We also include two counterfactual methods jointly estimating position bias and item relevance, the dual learning algorithm (DLA(Ai et al., 2018a) and pairwise debiasing (PairD(Hu et al., 2019). We introduce all unbiased learning to rank methods used in this work in Section 4.
2.2. ULTR on the Baidu-ULTR dataset
Most prior work on the Baidu-ULTR dataset has revolved around three public competitions: Two tracks at the WDSM Cup 2023 –unbiased learning-to-rank (Zou et al., 2023b) and pretraining for web search (Zou et al., 2023a) – and the ULTRE-2 track at NTCIR (Niu et al., 2023).
[…]
The second task of the WSDM Cup allowed participants to pre-train language models from clicks before fine-tuning on expert annotations. All top teams combined the output of their BERT models with traditional LTR features and tuned gradient-boosting models on annotations (Li et al., 2023b, a; Sun et al., 2023). Only the third-placed team incorporated a conventional ULTR approach with a softmax ranking loss and IPS during BERT pretraining.
In the NTCIR 17 ULTRE-2 track (Niu et al., 2023), the organizers released a subset of the data ($\approx 1$ million sessions) to make Baidu-ULTR more accessible, using the best-performing BERT model trained by Li et al. (2023b) during the WSDM Cup to create query-document embeddings in combination with traditional lexical matching features. [^6] Note that the first ULTRE-1 task at NTCIR 16 was not yet conducted on Baidu-ULTR but on a semi-synthetic simulation setup (Zhao et al., 2022). In their baseline experiments, they find that DLA models perform better than non-ULTR models trained with a pointwise loss but only marginally better than those trained with a listwise loss. The only participating team, Yu et al. (2023a), proposed to alleviate selection bias on items that are relevant but never clicked by using a DLA model to re-annotate non-clicked documents. However, they do not compare with other ULTR and non-ULTR baselines.
Overall, participants in the three cups reported noticeably higher ranking performance than the original experiments by the Baidu-ULTR dataset authors (Zou et al., 2022), and identified helpful techniques, such as re-weighting queries to address the long-tail query distribution (Sun et al., 2023), negative sampling of documents (Sun et al., 2023), ignoring clicks on items displayed for a short time (Li et al., 2023b), or pseudo-relevance feedback (Yu et al., 2023a). However, it remains unclear whether observed improvements stemmed from debiasing the click feedback or such orthogonal techniques. In order to identify the contribution of click debiasing to the results, we extend the work by Niu et al. (2023) and Zou et al. (2022) to larger datasets, more ULTR methods, fairer baselines, and investigate the interactions of ULTR with language model training. Next, we introduce the dataset used in this work in more detail.
3. Overview of Baidu-ULTR
3.1. Description of the data
The Baidu-ULTR dataset (Zou et al., 2022) contains more than $1.2$ billion search sessions (split into $2,000$ dataset partitions) with user clicks and $7,008$ annotated queries for evaluation. The dataset was collected in April 2022 by randomly sampling the search traffic of Baidu (Zou et al., 2022), reflecting the long-tail query distribution of real user traffic. The dataset contains logged queries, document titles, and abstracts of the search results presented to the user. The authors tokenized all released text with a private vocabulary for user privacy. In addition to tokenized text, the dataset also contains a multitude of logged user interactions, including clicks, dwell-time, and skipping behavior, as well as item presentation features, including document height, item type, and ranking position. In this work, we solely focus on item position and user clicks.
To foster the reproducibility of this work and to ease access to the vast Baidu-ULTR dataset for academic research, we release two smaller datasets: A language modeling dataset, comprised of the first 125 partitions of the dataset for training transformer models from scratch, and a smaller reranking dataset comprised of four partitions to compare ULTR on pre-computed query-document features. In the following, we describe the preprocessing common to both datasets and analyze the properties of our reranking dataset.
3.1.1. Pre-processing
(i) We use the md5 hash of the query tokens and document URL, respectively, as query and document identifiers. (ii) We discovered that over 20% of the Baidu-ULTR dataset comprised only two title-abstract token combinations. To avoid these documents polluting model training, we remove the what other people searched item from the ranking ($\approx 9\%$ of documents) and skip all documents with only a dash in the title, indicating no available content ($\approx 13\%$ of documents). [^7] While Baidu-ULTR was tokenized with a private vocabulary, Zou et al. (2022) confirmed to us that the tokens: $[3742,0111492,0112169,015061,0116905]$ translate to what other people searched and token $21429$ translates to “-” indicating missing content. Note that documents displayed after a skipped document keep their original position and do not, e.g., move to the position of a skipped document. (iii) Third, we drop all queries with less than five documents to display, affecting $\approx 2\%$ of the train queries and $\approx 0.3\%$ of annotated queries, leaving a remainder of 6,985 annotated test queries.
In contrast to previous work (Niu et al., 2023), we do not remove queries without any clicks. While this is common in ULTR as sessions without clicks do not contribute to popular pairwise ranking losses (Joachims et al., 2017), they are essential to train and evaluate methods predicting calibrated click probabilities.
3.1.2. Analysis of the reranking dataset
The reranking dataset comprises three dataset partitions for training ($\approx 1.8$ million user sessions, $\approx 11.7$ million query-document pairs) and one partition for validation and testing ($\approx 590k$ user sessions, $\approx 4.8$ million query-document pairs). Table 1 gives an overview of the statistics of the reranking dataset. While the dataset is a fraction of Baidu-ULTR, we highlight that it is still substantially larger than existing ULTR datasets (Ai et al., 2018b) and contains more unique queries and query-document pairs than the LTR datasets commonly used in semi-synthetic simulation for ULTR (Dato et al., 2022, 2016; Qin and Liu, 2013; Chapelle and Chang, 2011).
Besides query-document text and identifiers (as described in the preprocessing section), the reranking dataset also contains three sets of pre-computed query-document features: the CLS token of the BERT cross-encoder released by Zou et al. (2022), the CLS token of a BERT cross-encoder trained by us in the same fashion as the original model, and classic lexical learning-to-rank (LTR) features, including BM25 (Robertson et al., 1994), Tf-IDF (Robertson, 2004), and query-likelihood (with Jelinek-Mercer and Dirichlet smoothing (Chen and Goodman, 1999)). We compute the required inverted index for the methods above on the language modeling dataset that we use to train our BERT cross-encoder for fair comparison. As we restrict ourselves to solely training on click data and only use the expert annotations in the final evaluation, we compute all lexical matching features with default parameters instead of tuning them on annotations. Also, LTR parameters published by  Chen et al. (2023a) do not lead to major improvements over untuned default parameters in our setting. We publish two versions of the dataset on Huggingface, one using the Baidu cross-encoder [^8] https://huggingface.co/datasets/philipphager/baidu-ultr_baidu-mlm-ctr and one with our cross-encoder [^9] https://huggingface.co/datasets/philipphager/baidu-ultr_uva-mlm-ctr with a list of all pre-computed lexical ranking features and their respective hyperparameters.
As the reranking dataset is a random sample of the larger (preprocessed) dataset, we can use it to analyze the basic properties of Baidu-ULTR. After preprocessing, the average search session contains $8$ documents, with an average of $0.68$ clicks per session (number of clicks/number of sessions) and $\approx 46.5\%$ of sessions containing at least one click (see Table 1). Notably, we measure the overlap between queries (md5 hash of query tokens) in the click dataset and the annotated test set and find that only $12.7\%$ of annotated queries occur in the training dataset. While they appear during training, they make up $<0.1\%$ of the training dataset.
Table 1. Description of the reranking dataset we derive from Baidu-ULTR (Zou et al., 2022); statistics after preprocessing.
Implicit feedback
Annotations train test
Unique queries $6,985$ $1,378,901$ $501,215$
Unique documents $381,552$ $9,455,953$ $3,557,825$
Total query/doc pairs $382,038$ $11,715,447$ $4,209,900$
Total sessions - $1,779,017$ $593,930$
Total impressions - $14,526,276$ $4,848,878$
[…]
% of sessions with $\geqslant 1$ click $46.581\%$
% of sessions with $\geqslant 2$ clicks $13.082\%$
% of train queries occurring in the annotated set $\phantom{1}0.064\%$
% of annotated queries occurring in the train set $12.713\%$
3.2. Position bias
Standard counterfactual LTR methods using IPS require an estimation of position bias, which we estimate on our reranking dataset. We implement three intervention harvesting techniques that leverage the co-occurrence of the same query-document pair at different ranks (Agarwal et al., 2019b). Intervention harvesting mines query-document pairs logged in various positions, ideally due to distinct rankers in an A/B test ranking items differently. We can use this natural variability of encountering a document in different positions to estimate bias using click ratios between neighboring positions (Adjacent Pair), each position and a fixed rank (Pivot Rank), or between all ranking positions (All Pairs). We refer the reader to Agarwal et al. (2019b) for a detailed introduction to intervention harvesting.
We also estimate position bias using RegressionEM (REM) (Wang et al., 2018), which leverages query-document embeddings. REM estimates position bias over co-occurrences of query-document pairs with similar features rather than strictly identical query-document pairs, like intervention harvesting. In our case, we use the semantic embeddings of our naive BERT cross-encoder. Figure 2 displays our propensity estimations and the click-through rate per rank. The estimated position bias is broadly consistent across methods. REM and intervention harvesting arrive at similar estimations from different query-document representations. Our finding suggests the presence of a noticeable, top-heavy position bias affecting user behavior on Baidu-ULTR for which ULTR methods should be helpful.
Figure 2. Position bias as estimated by RegressionEM and three intervention harvesting methods compared to the mean CTR. Propensities were normalized by position one.
4. Unbiased Learning-to-Rank Methods
This section introduces the different ULTR methods used throughout this work. Similarly to Zou et al. (2022), we restrict ourselves to methods that address position bias according to the PBM and focus on benchmarking the learning algorithm used by each method.
First, we introduce the notation we will use to describe each method. Let $q$ be a user query and $D_{q}$ the ordered list of documents for $q$. Let $d\in D_{q}$ be a document in the list displayed at position $k$. We use $R_{q,d}$, $E_{k}$ and $C_{q,d,k}$, respectively, to denote the random events for: (i) document $d$ being relevant for query $q$, (ii) position $k$ of the result page being examined by the user, and (iii) document $d$ placed in position $k$ for query $q$ being clicked. We assume users follow the position-based model (PBM), meaning that users click only if they observed the position $k$ of an item and deemed the displayed document $d$ to be relevant:
(1) $P(C_{q,d,k})=P(E_{k})\times P(R_{q,d}).$
In the following, we use $\mathcal{L}$ to reference a ranking loss function; in this work, either binary cross-entropy, softmax cross-entropy (Bruch et al., 2019), or LambdaRank (Burges et al., 2006). All ranking methods aim to estimate document relevance $r(q,d)=P(R_{q,d})$ and optionally examination $e(k)=P(E_{k})$ (or related quantities) from observed clicks $c\in\{0,1\}$. We use $\tilde{x}$ to denote estimators of a quantity $x$. We can now define the methods compared in our study:
Naive: is the common ULTR term for applying a ranking loss without any position bias correction. This means that we naively interpret clicks as positive/negative user feedback:
(2) $\mathcal{L}^{\textup{naive}}=\mathcal{L}(\tilde{c}(q,d);c).$
We implement a pointwise version of this model using binary cross-entropy and two listwise versions using softmax cross-entropy (Bruch et al., 2019) and LambdaRank (Burges et al., 2006), respectively.
Two-Tower (Yan et al., 2022, Two-Tower Model): jointly learns the maximum-likelihood parameters of a relevance model $\tilde{r}$ and an examination model $\tilde{e}$, directly mirroring the PBM assumptions:
(3) $\begin{split}\mathcal{L}^{\textup{TT}}=\mathcal{L}\left(\sigma\left(\tilde{e}(k)+\tilde{r}(q,d)\right);c\right).\end{split}$
We use the additive formulation (Yan et al., 2022; Zhang et al., 2023) with $\tilde{r}$ and $\tilde{e}$ representing logits, $\sigma$ the sigmoid function, and $\mathcal{L}$ the binary cross-entropy loss.
RegressionEM (Wang et al., 2018, Regression expectation-maximization): learns a maximum likelihood estimate of relevance and examination parameters through expectation-maximization (Dempster et al., 1977):
(4) $\begin{split}\mathcal{L}^{\textup{REM}}={}&\mathcal{L}(\tilde{r}(q,d);c+(1-c)\overline{r}(q,d))\\ &+\mathcal{L}\left(\tilde{e}(k);c+(1-c)\overline{e}(k)\right),\end{split}$
where $\overline{r}(q,d)=\frac{\tilde{r}(q,d)\left(1-\tilde{e}(k)\right)}{1-\tilde{r}(q,d)\tilde{e}(k)}$ and $\overline{e}(k)=\frac{\tilde{e}(k)\left(1-\tilde{r}(q,d)\right)}{1-\tilde{r}(q,d)\tilde{e}(k)}$ are the posterior relevance and examination probabilities. In practice, we use binary cross-entropy as the base loss and replace the original expectation-maximization procedure with a numerically stable gradient-based version suggested in TF-Rank (Pasumarthi et al., 2019) by performing the posterior computation with logits.
IPS (Joachims et al., 2017, Inverse propensity scoring): Given known examination probabilities $e_{k}$, IPS re-weights the click labels by the propensity (i.e., normalized examination probability) of the current position:
(5) $\mathcal{L}^{\textup{IPS}}=\mathcal{L}\left(\tilde{r}(q,d);\frac{\text{max}(\tau,e_{1})}{\text{max}(\tau,e_{k})}c\right).$
In our experiments, we use propensities estimated by the AllPairs intervention harvesting method (Agarwal et al., 2019b) and reduce variance by clipping examination probabilities to a minimum value of $\tau=0.1$ (Joachims et al., 2017). We implement a pointwise (Bekker et al., 2019; Saito et al., 2020) and listwise IPS variant (Bruch et al., 2019; Ai et al., 2018a).
DLA (Ai et al., 2018a, Dual learning algorithm): separately learns tunable relevance scores with fixed propensities and tunable propensities with fixed relevance scores, essentially applying IPS twice:
(6) $\mathcal{L}^{\textup{DLA}}=\mathcal{L}\left(\tilde{r}(q,d);\frac{\tilde{e}(1)}{\tilde{e}(k)}c\right)+\mathcal{L}\left(\tilde{e}(k);\frac{\tilde{r}(q,d_{1})}{\tilde{r}(q,d)}c\right),$
where $d_{1}$ is the document ranked first on the current result page. We use softmax cross-entropy as the base loss and perform a softmax-normalization of examination and relevance probabilities $\tilde{r}$ and $\tilde{e}$ within a result page, as in the original paper (Ai et al., 2018a).
PairD (Hu et al., 2019, Pairwise debiasing): learns positive $\tilde{e}^{+}$ and negative $\tilde{e}^{-}$ propensities (i.e., corresponding to clicks and non-clicks, respectively) with a constraint on the norm of the learned propensities:
(7) $\mathcal{L}^{\textup{PairD}}=\frac{\mathcal{L}(\tilde{r}(q,d);c)}{\tilde{e}^{+}(k)\tilde{e}^{-}(k)}+\left\|\tilde{e}^{+}(k)\right\|+\left\|\tilde{e}^{-}(k)\right\|.$
In practice, we use the $L_{1}$-norm for propensity regularization and learn the relevance scores using the LambdaRank loss (Burges et al., 2006), as in the original paper (Hu et al., 2019). Note that the theoretical validity of this method under non-trivial position bias has been challenged by Oosterhuis (2022). Still, we include it in our comparison as it has demonstrated strong empirical performance in past comparisons (Ai et al., 2021).
5. Experimental Setup
5.1. Training and evaluation procedure
All models are trained on our reranking dataset, and we use a 50/50 random split of our test click dataset (see Table 1) for validation and testing. We evaluate ranking performance on $6,985$ annotated queries, where experts rated each query-document pair’s relevance on a scale from 00 to $4$. We point to Zou et al. (2022) for more details on the annotation process. To remain focused, we do not use available side information such as dwell-time, off-screen scrolling, and returns to the result page. Instead, we use query and document content, document position and click label as our training data.
We measure ranking performance on the annotated set using discounted cumulative gain (DCG) at different truncation levels and mean reciprocal rank (MRR) at $10$, as well as negative log-likelihood (NLL) for click prediction.
5.2. Models
Traditionally, unbiased learning-to-rank practitioners train small ranking models with LTR features such as BM25 (Robertson et al., 1994) or TF-IDF (Robertson, 2004) as input (Ai et al., 2018a; Oosterhuis and de Rijke, 2021; Vardasbi et al., 2020b). To investigate the role of semantic embeddings and the interaction of ULTR with language model training on such a large dataset, we train two types of models: transformer-based language models trained from scratch following the MonoBERT cross-encoder (Nogueira et al., 2019), and multi-layer perceptron-based reranking models, that take as input either traditional LTR features or fixed query-document embeddings obtained by the language models.
5.2.1. Language models
We train cross-encoders on the Baidu-ULTR dataset from scratch in Jax (Bradbury et al., 2018), building on the FlaxBERT implementation from Huggingface. [^10] https://huggingface.co/docs/transformers/model_doc/bert#transformers.FlaxBertForPreTraining The input to our BERT models is: [CLS] query [SEP] doc_title [SEP] doc_abstract. Following previous work on Baidu-ULTR (Zou et al., 2022; Chen et al., 2023b), we truncate the input to a maximum length of 128 tokens. Compared to the original BERT training scheme (Devlin et al., 2019), we keep the masked-language modeling task (masking $30\%$ of input tokens at random) but discard the next-sentence prediction task. Instead, we follow the MonoBERT (Nogueira et al., 2019) setup and apply a linear layer to the BERT CLS token to output a click prediction score. We train multiple models with different click-based loss functions: naive click prediction with a pointwise (binary cross-entropy) and a listwise (softmax cross-entropy) loss, pointwise and listwise IPS, a pointwise two-tower loss, and a listwise DLA loss (see Section 4). All loss functions were built on top of the Rax library (Jagerman et al., 2022).
We keep the architecture of the original $\text{BERT}_{\text{base}}$ model, using 12 transformer layers, 12 attention heads, 768 output dimensions, and, following Chen et al. (2023b), a vocabulary size of $22,000$. All language models are trained with a batch size of $256$ ($4\times 64$) for $2$ million gradient steps, i.e., on $512$ million documents. Training each base model on four GPUs (NVIDIA H100-80GB) takes around 46 hours, which is a speed-up of almost 50% compared to an early PyTorch implementation of ours using the exact same library, architecture, and hardware. We use the AdamW optimizer (Loshchilov and Hutter, 2019) with a fixed learning rate of $5\times 10^{-5}$ and a weight decay of $0.01$. Given the substantial compute invested to train each language model, we rely on prevalent default parameters for BERT (listed in our repository).
5.2.2. Reranking models
[…]
5.3. Hyperparameter tuning
We perform extensive hyperparameter tuning for a fair comparison of the reranking models in our experiments. Given the immense combinatorial space of hyperparameters, methods, and datasets, we adopt an incremental hyperparameter tuning strategy as advocated by Godbole et al. (2023). First, we tune our model architecture per dataset based on the pointwise naive model. We tune the number of hidden dimensions $\in\{64,128,256,512,1024\}$ and the number of layers $\in\{2,3,4,5\}$ over three random seeds. As both network depth and width can interact with the learning rate, we tune each parameter combination over three learning rates $\in\{0.001,0.0005,0.0001\}$, i.e., we treat the learning rate as a nuisance parameter (Godbole et al., 2023). We adopt this incremental tuning procedure as we found no major discrepancies in model architecture between ULTR methods but instead between datasets. Subsequently, we adopt the architecture of the pointwise naive model per dataset for all other methods.
A five-layer perceptron with 512 hidden dimensions yields optimal results for our LTR features, while a five-layer perceptron with 256 dimensions performs best for the Baidu BERT embeddings. Additionally, a two-layer perceptron with 256 dimensions was the most suitable choice for our BERT embeddings. Given each base model architecture, we tune the final dropout $\in\{0,0.3\}$ and learning rate $\in\{0.001,0.0005,0.0003,0.0001\}$ for each method and dataset over three random seeds. We use the AdamW optimizer (Loshchilov and Hutter, 2019) with $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\varepsilon=1e^{-8}$. Most models work best with a learning rate of $0.0001$ and do not benefit from dropout. The full list of hyperparameters is available in our online repository. Given the final hyperparameters for each method, we train all methods for 50 epochs, stopping early after five epochs of no improvement of the validation loss computed on clicks. Lastly, we highlight that while this tuning step improved the performance of all methods, our main findings are consistent across many different hyperparameter combinations.
6. Results
Figure 3. Comparing ULTR method on pre-trained BERT embeddings and LTR features. We display the average ranking performance measured in DCG@10 over five independent runs and plot a bootstrapped 95% confidence interval. The grey dotted line indicates the performance of a random ranker.
Before presenting our findings, we recall our research questions:
  1. (RQ1)
    Does unbiased learning-to-rank improve performance on the Baidu-ULTR dataset over naive, non-debiasing models?
  2. (RQ2)
    How do ULTR methods fare against each other, and how do ranking losses and input features affect their performance?
  3. (RQ3)
    Can ULTR methods be applied during language model training and do they bring improved performance?
6.1. Unbiased learning-to-rank yields no or tiny improvements on Baidu-ULTR
In our first experiment, we train reranking models with pointwise and listwise ULTR loss functions and their respective naive, non-debiasing counterparts on three different types of input features: the CLS token of Zou et al. (2022)’s pointwise cross-encoder, the CLS token of our pointwise cross-encoder, and learning-to-rank features (see Section 3.1.2). We list comprehensive results in Table 3. First, we focus on the ranking performance on expert annotations as measured in DCG and MRR. We display the DCG@10 in Figure 3.
At first glance at Figure 3, echoing previous work (Zhao et al., 2022; Niu et al., 2023), we observe that we can get a much higher ranking performance from our cross-encoder than the initially released BERT cross-encoder. Models trained on LTR features and even of a trivial baseline assigning random scores (grey dotted line) to the annotated documents beat all models trained on the original Baidu BERT features.
Second, given a set of input features, the choice of ranking loss used as a base for ULTR methods matters greatly (comparing the colored groups in Figure 3). Methods based on LambdaRank outperform methods based on the listwise softmax loss, which outperform those based on the pointwise binary cross-entropy loss. In contrast, applying ULTR techniques yield marginal improvements at most compared to their naive, non-debiasing counterpart.
In detail, we observe that the pointwise two-tower and IPS models do not consistently and significantly improve performance compared to the naive pointwise loss – IPS is even significantly worse on both BERT features – and RegressionEM is the only pointwise ULTR method that brings small but sometimes significant improvements. Regarding listwise methods, IPS and DLA significantly but marginally improve on our BERT and LTR features. Finally, PairD is no better and sometimes worse than its naive LambdaRank counterpart. These results are broadly consistent across input features, suggesting that standard ULTR methods struggle to bring improvement to Baidu-ULTR, regardless of the query-document features.
Overall, we can answer research questions (RQ1) and (RQ2): on the Baidu-ULTR dataset, unbiased learning-to-rank techniques do not bring any clear benefits, particularly when contrasted with the significant and reliable differences based on the choice of query-document features and ranking loss. Our reranking results confirm the initial findings of Zou et al. (2022).
6.2. Language model training is sensitive to the choice of unbiased learning-to-rank method
Given the poor performance of ULTR on our reranking datasets (see Section 6.1), we further investigate whether training language models with ULTR is a promising direction for future research. We train six MonoBERT (Nogueira et al., 2019) cross-encoders with different pointwise and listwise loss functions, including ULTR loss functions.
Table 2 gives an overview of applying ULTR methods during language model training. In contrast to the inefficacy of ULTR-based rerankers, we observe stark differences due to applying ULTR methods when training language models. Similar to the reranking task, approaches based on the listwise softmax loss perform better than those based on pointwise binary cross-entropy. The additional application of ULTR brings considerable improvements with the pointwise two-tower objective. However, we also observe substantial degradations with both IPS and DLA. The best method is the naive listwise softmax loss without any debiasing objective.
These inconclusive results make us cautious about answering (RQ3): ULTR methods substantially impact language model training more than reranking with fixed embeddings. However, the influence of ULTR on ranking performance and learned query-document representations needs further investigation.
6.3. Click prediction does not imply ranking performance on annotations
In addition to evaluating ranking performance, we display the negative log-likelihood (NLL) of predicted click scores in Figure 4. We restrict this evaluation to pointwise ULTR methods, making well-defined click predictions. All ULTR methods robustly improve click prediction compared to a naive loss, with the two-tower model showing the largest improvement on all three datasets. Moreover, as discussed in Section 3.2, the propensities learned by RegressionEM are close to those discovered through intervention harvesting and show a strong position bias. This finding suggests that unbiased learning-to-rank robustly captures position bias and can better predict user clicks because of it. Yet, this improved click prediction does not translate to enhanced ranking performance on expert annotations.
Critically, we can see in Table 3 that ranking documents according to their (untuned) BM25 (Robertson et al., 1994) scores yields better ranking performance than all click-based models we trained. Even more surprisingly, using BM25 scores as part of the input for the reranking models – it is one of the LTR features – and training these models on clicks lowers ranking performance. This phenomenon was also described by Sun et al. (2023) during the WSDM Cup. These concurrent results suggest that click prediction and ranking performance on annotations are diverging objectives on Baidu-ULTR and training models on clicks does not guarantee improvements in ranking metrics.
Figure 4. Click prediction performance of pointwise methods measured in negative log-likelihood, lower is better.
7. Discussion
In light of the puzzling results reported in Section 6, we review potential reasons for the failure of ULTR in Section 7.1 as well as the limitations of our study in Section 7.2. Finally, we discuss the implications for the field of unbiased learning-to-rank in Section 7.3.
Table 2. Comparison of cross-encoder models trained from scratch on the Baidu-ULTR dataset with ULTR loss functions.
Model DCG@10 $\uparrow$ NLL $\downarrow$
Pointwise Naive $7.251$ $0.227$
Pointwise Two-Tower $7.456$ $0.217$
Pointwise IPS $6.296$ $0.317$
Listwise Naive $8.478$
Listwise IPS $7.450$
Listwise DLA $7.802$
7.1. Potential reasons for the failure of ULTR
No position bias. A first straightforward explanation for the inefficacy of ULTR techniques would be that position bias is not prevalent in the Baidu-ULTR dataset. However, our results in Section 3.2 show a strongly decreasing click-through rate with increasing position and a strong position bias estimated by techniques based on different methodologies. Recall that intervention harvesting uses CTR ratios of query-document pairs appearing in multiple positions while RegressionEM uses our BERT features and expectation maximization. The fact that these results agree strongly suggests the existence of position bias in this dataset and that ULTR methods like RegressionEM and Two-Tower models could capture it.
More complex user behavior. Another hypothesis, initially suggested by Zou et al. (2022), is that the actual user model is more complicated than the simple PBM we consider in this work. This possibility might explain why listwise approaches bring such considerable improvements over pointwise methods and would not be surprising as user studies have identified more complex biases (Agarwal et al., 2019a; Zhuang et al., 2021; Sarvi et al., 2023). However, we point out that ULTR methods bringing no benefits would mean that the substantial position bias we identified is negligible against other types of biases. This assumption, in turn, would contradict related work where PBM-based models trained on more complex user behaviors still improved performance, albeit by not as much as the correct user model (Vardasbi et al., 2020a; Deffayet et al., 2023b).
Lack of identifiability. Past studies have identified that a lack of variability in the logged data, i.e., not encountering a query-document pair at different positions, can lead to the dataset not being identifiable, in the sense that there exist infinite valid combinations of relevance and bias parameters (Oosterhuis, 2022; Chen et al., 2024). However, as this work’s four bias estimation methods converged to a similar position bias, we deem this explanation rather unlikely.
Distribution shift. Poor ranking performance may also be caused by the distribution shift between the query-document pairs in the training and annotated test set. While methods train only on the top-$10$ search results, the annotation dataset also includes presumably much less relevant documents sampled from the top-$1000$ candidate documents. This shift in the input feature distribution, in conjunction with the low overlap between train and test queries (see Section 3.1.2) and the long-tail query distribution during training, might all potentially impact ULTR. More research and, ideally, clicks for queries with expert annotations are needed to better understand the impact of the various distribution shifts on ULTR.
Strong logging policy. While we do not have access to logging policy scores and rankings on the annotated test queries, the limitations imposed by an already strong production system have been well-documented (Deffayet et al., 2023a). A side observation of ours reinforces this possibility: methods that show strong ranking performance correlate highly with the logging policy of the click dataset (as estimated through the original item position). For future dataset releases, we recommend including logging policy scores to enable assessments of the improvement made by training new rankers on click data compared to the logging policy (Deffayet et al., 2023a; Gupta et al., 2023).
User-annotator disagreement. Finally, the divergence of click prediction and ranking performance hints at a deeper issue: annotators may not find the same documents relevant as users scrolling through a result page, which might be personalized and part of a multi-query session to find actionable information for the current user needs. We hope more user studies or parallel datasets of clicks and expert annotations can clarify this hypothesis.
[…]
Input features Method DCG@1 $\uparrow$ DCG@3 $\uparrow$ DCG@5 $\uparrow$ DCG@10 $\uparrow$ MRR@10 $\uparrow$ NLL $\downarrow$
Random $1.472$ $(0.020)$ $3.148$ $(0.030)$ $4.349$ $(0.036)$ $6.693$ $(0.044)$ $0.577$ $(0.003)$ $0.944$ $(0.002)$
BM25 $(k_{1}=1.2,b=0.75)$ $2.211$ $4.656$ $6.377$ $9.544$ $0.716$
Baidu BERT Pointwise Naive $1.264$ $(0.030)$ $2.765$ $(0.059)$ $3.881$ $(0.068)$ $6.111$ $(0.079)$ $0.535$ $(0.007)$ $0.246$ $(0.005)$
Pointwise Two-Tower $1.266$ $(0.018)$ $2.769$ $(0.023)$ $3.893$ $(0.026)$ $6.114$ $(0.031)$ $0.533$ $(0.003)$ $0.228$ $(0.005)$
Pointwise RegressionEM $1.343$ $(0.034)$ $2.884$ $(0.074)$ $4.017$ $(0.091)$ $6.257$ $(0.107)$ $0.548$ $(0.005)$ $0.229$ $(0.005)$
Pointwise IPS $1.157$ $(0.011)$ $2.606$ $(0.012)$ $3.699$ $(0.013)$ $5.915$ $(0.014)$ $0.517$ $(0.002)$ $0.230$ $(0.005)$
Listwise Naive $1.362$ $(0.022)$ $2.940$ $(0.037)$ $4.096$ $(0.037)$ $6.388$ $(0.042)$ $0.549$ $(0.004)$
Listwise IPS $1.353$ $(0.034)$ $2.920$ $(0.060)$ $4.078$ $(0.063)$ $6.359$ $(0.071)$ $0.548$ $(0.005)$
Listwise DLA $1.345$ $(0.020)$ $2.917$ $(0.037)$ $4.083$ $(0.048)$ $6.378$ $(0.051)$ $0.549$ $(0.004)$
LambdaRank Naive $1.419$ $(0.017)$ $3.047$ $(0.039)$ $4.237$ $(0.054)$ $6.570$ $(0.065)$ $0.562$ $(0.005)$
LambdaRank PairD $1.410$ $(0.028)$ $3.025$ $(0.044)$ $4.215$ $(0.064)$ $6.550$ $(0.066)$ $0.559$ $(0.005)$
Our BERT Pointwise Naive $1.705$ $(0.013)$ $3.602$ $(0.022)$ $4.944$ $(0.023)$ $7.546$ $(0.031)$ $0.619$ $(0.002)$ $0.221$ $(0.005)$
Pointwise Two-Tower $1.656$ $(0.017)$ $3.556$ $(0.024)$ $4.948$ $(0.032)$ $7.639$ $(0.038)$ $0.615$ $(0.004)$ $0.213$ $(0.006)$
Pointwise RegressionEM $1.694$ $(0.031)$ $3.619$ $(0.056)$ $4.998$ $(0.081)$ $7.657$ $(0.111)$ $0.618$ $(0.004)$ $0.215$ $(0.005)$
Pointwise IPS $1.589$ $(0.012)$ $3.438$ $(0.012)$ $4.794$ $(0.013)$ $7.436$ $(0.035)$ $0.604$ $(0.001)$ $0.214$ $(0.005)$
Listwise Naive $1.798$ $(0.008)$ $3.768$ $(0.020)$ $5.167$ $(0.029)$ $7.844$ $(0.046)$ $0.631$ $(0.001)$
Listwise IPS $1.816$ $(0.017)$ $3.800$ $(0.027)$ $5.200$ $(0.031)$ $7.885$ $(0.043)$ $0.634$ $(0.002)$
Listwise DLA $1.816$ $(0.011)$ $3.806$ $(0.022)$ $5.210$ $(0.024)$ $7.890$ $(0.031)$ $0.634$ $(0.001)$
LambdaRank Naive $1.931$ $(0.016)$ $3.993$ $(0.029)$ $5.452$ $(0.040)$ $8.233$ $(0.055)$ $0.648$ $(0.003)$
LambdaRank PairD $1.932$ $(0.007)$ $3.985$ $(0.019)$ $5.438$ $(0.022)$ $8.216$ $(0.038)$ $0.646$ $(0.003)$
LTR Features Pointwise Naive $1.312$ $(0.034)$ $2.977$ $(0.066)$ $4.246$ $(0.090)$ $6.757$ $(0.133)$ $0.543$ $(0.008)$ $0.247$ $(0.005)$
Pointwise Two-Tower $1.333$ $(0.046)$ $2.986$ $(0.085)$ $4.241$ $(0.094)$ $6.698$ $(0.110)$ $0.553$ $(0.008)$ $0.228$ $(0.006)$
Pointwise RegressionEM $1.397$ $(0.068)$ $3.062$ $(0.133)$ $4.300$ $(0.166)$ $6.729$ $(0.181)$ $0.559$ $(0.014)$ $0.229$ $(0.005)$
Pointwise IPS $1.352$ $(0.044)$ $3.013$ $(0.095)$ $4.258$ $(0.126)$ $6.717$ $(0.154)$ $0.560$ $(0.011)$ $0.232$ $(0.005)$
Listwise Naive $1.599$ $(0.023)$ $3.485$ $(0.042)$ $4.845$ $(0.072)$ $7.443$ $(0.100)$ $0.596$ $(0.005)$
Listwise IPS $1.641$ $(0.020)$ $3.545$ $(0.012)$ $4.920$ $(0.020)$ $7.522$ $(0.026)$ $0.602$ $(0.002)$
Listwise DLA $1.621$ $(0.026)$ $3.525$ $(0.074)$ $4.891$ $(0.101)$ $7.512$ $(0.160)$ $0.599$ $(0.006)$
LambdaRank Naive $1.750$ $(0.035)$ $3.717$ $(0.068)$ $5.125$ $(0.083)$ $7.810$ $(0.111)$ $0.613$ $(0.007)$
LambdaRank PairD $1.723$ $(0.018)$ $3.683$ $(0.034)$ $5.089$ $(0.042)$ $7.761$ $(0.046)$ $0.608$ $(0.004)$
7.2. Limitations
First, we only considered the correction of position bias under the position-based model, while other biases might have to be mitigated (Agarwal et al., 2019a; Vardasbi et al., 2020a). Second, we did not include bias features available in Baidu-ULTR beyond positions. We highlight that Chen et al. (2023b) report improved ranking performance by considering additional bias features, including item height or media type. Third, the variability in item position that we used for position bias estimation is likely not due to natural variability but due to unobserved confounding and algorithmic choices based on additional context information rather than a stochastic logging policy. Fourth, our work only applied ULTR during language model pre-training and on fixed BERT embeddings for reranking. We did not yet explore multi-stage pre-training schemes, such as pre-training BERT on a naive click prediction task and later fine-tuning BERT using ULTR.
Lastly, we stress that our findings are strictly limited to the Baidu-ULTR dataset. While the dataset was collected from one of the most used search engines globally, it cannot be considered representative of all real-world search scenarios. Yet, we believe that results on this large-scale dataset are important for the research community.
7.3. Implications for the field
Our results starkly contrast with many studies in semi-synthetic simulation (Joachims et al., 2017; Ai et al., 2021; Tran et al., 2021; Saito et al., 2020; Ai et al., 2018a; Hu et al., 2019). They call for adjusting these semi-synthetic experimental setups to better reflect real-world challenges. The richness of this dataset allows for seeding simulations with more plausible data and, therefore, might bridge the gap between simulations and reality. Moreover, the prevalence of transformer-based models and text embeddings in current IR research encourages exploring the interaction of ULTR methods with such models. As shown in Section 6.2, the same techniques applied during language model training, instead of on a small re-ranking model, can yield vastly different results.
More broadly, the puzzling results we obtained, especially the apparent divergence between clicks and relevance annotations, prompt us to rethink how we measure success in ULTR. Expert annotations are static and might not reflect user context, so results obtained on annotated datasets can be polluted by distribution shifts or logging policy performance. In particular, we believe that, whenever possible, we should evaluate ULTR methods on the tasks they are trained to accomplish, including relevance estimation, bias estimation, CTR maximization, fairness of exposure, etc.
Finally, we would like to stress that this paper is not a judgment on the ULTR field. As mentioned above, results may differ on other datasets, and most ULTR methods are theoretically justified, meaning only their validity in this scenario has been challenged. In fact, our experiments validate some intuitions formulated in the ULTR community. For instance, the stark differences between loss functions justify the motivation of Joachims et al. (2017) to connect simple user models with more powerful ranking loss functions.
8. Conclusion
In this work, we carefully revisited and extended the experiments conducted by Zou et al. (2022) on the recently released Baidu-ULTR dataset. As the largest publicly available dataset comprising both click logs from a major search engine and expert annotations, this dataset constitutes a rare opportunity to assess the progress of click-based learning-to-rank, and especially unbiased learning-to-rank.
Our main conclusion, however, is that while we have observed substantial improvements relating to the choice of query-document representations and ranking loss (e.g., pointwise or listwise), conventional unbiased learning-to-rank techniques do not bring any clear benefits when applied to the Baidu-ULTR dataset. Our findings confirm the work by Zou et al. (2022), even though we used different dataset features and preprocessing, model implementations, and performed extensive hyperparameter tuning. Further, we observed a divergence between click-based and annotations-based objectives, as all click-based approaches were outperformed on annotation-based metrics by simple baselines such as BM25, even when BM25 was a model input. We believe these results call for more research into the conditions for success and failure of unbiased learning-to-rank and click-based approaches as a whole.
Acknowledgements.
We thank Lixin Zou for his valuable insights and clarifications on the original Baidu work, as well as Jiaxin Mao and Zechun Niu for sharing their learnings from NTCIR 17. This research was supported by the Mercury Machine Learning Lab, created by TU Delft, the University of Amsterdam, and Booking.com. Maarten de Rijke was supported by the Hybrid Intelligence Center, a 10-year program funded by the Dutch Ministry of Education, Culture and Science through the Netherlands Organisation for Scientific Research, https://hybrid-intelligence-centre.nl. All content represents the opinion of the authors, which is not necessarily shared or endorsed by their respective employers and/or sponsors.

### v2

@@
 While theoretically justified and extensively tested in simulation, ULTR techniques lack empirical validation, especially on modern search engines.
-The dataset released for the WSDM Cup 2023, collected from Baidu’s search engine, offers a rare opportunity to assess the real-world performance of prominent ULTR techniques.
+The Baidu-ULTR dataset released for the WSDM Cup 2023, collected from Baidu’s search engine, offers a rare opportunity to assess the real-world performance of prominent ULTR techniques.
 Despite multiple submissions during the WSDM Cup 2023 and the subsequent NTCIR ULTRE-2 task, it remains unclear whether the observed improvements stem from applying ULTR or other learning techniques.
-We revisit and extend the available experiments.
-We find that unbiased learning-to-rank techniques do not bring clear performance improvements, especially compared to the stark differences brought by the choice of ranking loss and query-document features.
-Our experiments reveal that ULTR robustly improves click prediction.
-However, these gains in click prediction do not translate to enhanced ranking performance on expert relevance annotations, implying that conclusions strongly depend on how success is measured in this benchmark.
-submissionid: 4266journalyear: 2024copyright: rightsretainedconference: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 14–18, 2024; Washington D.C., USAbooktitle: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’24), July 14–18, 2024, Washington D.C., USA
+In this work, we revisit and extend the available experiments on the Baidu-ULTR dataset.
+We find that standard unbiased learning-to-rank techniques robustly improve click predictions but struggle to consistently improve ranking performance, especially considering the stark differences obtained by choice of ranking loss and query-document features.
+Our experiments reveal that gains in click prediction do not necessarily translate to enhanced ranking performance on expert relevance annotations, implying that conclusions strongly depend on how success is measured in this benchmark.
+Learning to rank, Counterfactual learning-to-rank, Click models
+journalyear: 2024copyright: rightsretainedconference: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 14–18, 2024; Washington, DC, USAbooktitle: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’24), July 14–18, 2024, Washington, DC, USAdoi: 10.1145/3626772.3657892isbn: 979-8-4007-0431-4/24/07
 1. Introduction
@@
 However, clicks are usually a biased signal of relevance, as they often depend on the position of an item (Joachims et al., 2005), its surrounding items (Craswell et al., 2008; Zhuang et al., 2021), or even the user’s trust in the system to place relevant items on top (Joachims et al., 2005; Agarwal et al., 2019a; Vardasbi et al., 2020b).
-Over the years, the field of unbiased learning-to-rank (ULTR) has proposed methods to mitigate biases when training ranking models on click data.
+Over the years, the field of unbiased learning-to-rank (ULTR) has proposed methods to mitigate biases when training ranking models on click data (Gupta et al., 2024).
 ULTR methods, especially in academic research, are often evaluated in synthetic or semi-synthetic setups (Joachims et al., 2017; Ai et al., 2018a; Vardasbi et al., 2020b, a; Oosterhuis and de Rijke, 2021).
@@
 Figure 1.
-Original comparison on Baidu-ULTR, showing four ULTR methods not significantly outperforming a naive method not correcting for position bias.
+The original experiments on Baidu-ULTR show that none of the four compared ULTR methods outperform a naive method not correcting for position bias.
 Data from (Zou et al., 2022, Table 6) and visualization by us.
@@
 Interestingly, Zou et al.
-(2022) find that these prominent ULTR methods do not perform significantly better than a naive BERT model trained on clicks without any bias correction, [^1] Zou et al.
-(2022) do find ULTR to be beneficial on frequent head queries compared to infrequent tail queries.
+(2022) find that these prominent ULTR methods overall do not perform significantly better than a naive BERT model trained on clicks without any bias correction, [^1] Zou et al.
+(2022) find that DLA outperforms a naive baseline on frequent head queries.
 as we report in Figure 1.
@@
   2. (2)
-    Three competitions have been conducted on the Baidu-ULTR dataset, with participants reporting vastly improved ranking results (although not necessarily using ULTR).
-A preliminary experiment by us revealed that a random ranking of the annotated test dataset achieves a DCG@10 $\approx 6.69$, substantially higher than all baselines in Figure 1 and thus, hinting at potential problems in the original experiments.
+    Three competitions have been conducted on the Baidu-ULTR dataset, with participants reporting vastly improved ranking results, though not necessarily using ULTR.
+A preliminary experiment by us revealed that a random ranking of the annotated test dataset achieves a DCG@10 $\approx 6.69$, which is higher than all baselines in Figure 1 and suggests potential issues with the original experiments.
   3. (3)
@@
 IPS, for example, requires an estimate of position bias on the given dataset, while the creators of the Baidu-ULTR dataset confirmed using a hardcoded position bias from the ULTRA library in their experiments (Tran et al., 2021).
-We also found that placeholder documents make up for almost $20\%$ of the Baidu-ULTR dataset (see Section 3), with an unknown impact on the trained models.
+We also found that placeholder documents make up almost $20\%$ of the Baidu-ULTR dataset (see Section 3), with an unknown impact on the trained models.
 By revisiting the experiments reported by Zou et al. (2022) and the cups’ participants, we address the following research questions:
@@
     Can ULTR methods be applied during language model training, and do they bring improved performance?
-We answer (RQ1) negatively as we cannot find robust and significant improvements in a fair comparison between ULTR-based and naive methods.
-We address (RQ2) by reporting minor but existing differences between ULTR methods and by showcasing large and reliable differences due to ranking losses and input features.
+We answer (RQ1) negatively as we cannot find robust improvements in ranking performance in a fair comparison between ULTR-based and naive methods.
+For (RQ2), we report minor but existing differences between ULTR methods and showcase large and reliable differences due to ranking losses and input features.
 Regarding (RQ3), we find that while certain ULTR methods improve over their naive counterpart if applied during language model training, the interactions of ULTR and transformer-based models are not well understood and can even lead to decreased performance.
@@
 Two-Tower (Yan et al., 2022, Two-Tower Model): jointly learns the maximum-likelihood parameters of a relevance model $\tilde{r}$ and an examination model $\tilde{e}$, directly mirroring the PBM assumptions:
-(3) $\begin{split}\mathcal{L}^{\textup{TT}}=\mathcal{L}\left(\sigma\left(\tilde{e}(k)+\tilde{r}(q,d)\right);c\right).\end{split}$
+(3) $\begin{split}\mathcal{L}^{\textup{TwoTower}}=\mathcal{L}\left(\sigma\left(\tilde{e}(k)+\tilde{r}(q,d)\right);c\right).\end{split}$
 We use the additive formulation (Yan et al., 2022; Zhang et al., 2023) with $\tilde{r}$ and $\tilde{e}$ representing logits, $\sigma$ the sigmoid function, and $\mathcal{L}$ the binary cross-entropy loss.
@@
 PairD (Hu et al., 2019, Pairwise debiasing): learns positive $\tilde{e}^{+}$ and negative $\tilde{e}^{-}$ propensities (i.e., corresponding to clicks and non-clicks, respectively) with a constraint on the norm of the learned propensities:
-(7) $\mathcal{L}^{\textup{PairD}}=\frac{\mathcal{L}(\tilde{r}(q,d);c)}{\tilde{e}^{+}(k)\tilde{e}^{-}(k)}+\left\|\tilde{e}^{+}(k)\right\|+\left\|\tilde{e}^{-}(k)\right\|.$
+(7) $\mathcal{L}^{\textup{PairD}}=\frac{\mathcal{L}(\tilde{r}(q,d);c)}{\tilde{e}^{+}(k)\tilde{e}^{-}(k)}+\|\tilde{e}^{+}\|_{1}+\|\tilde{e}^{-}\|_{1}.$
 In practice, we use the $L_{1}$-norm for propensity regularization and learn the relevance scores using the LambdaRank loss (Burges et al., 2006), as in the original paper (Hu et al., 2019).
@@
 (2022) for more details on the annotation process.
-To remain focused, we do not use available side information such as dwell-time, off-screen scrolling, and returns to the result page.
-Instead, we use query and document content, document position and click label as our training data.
+In this work, we do not use available side information such as dwell-time, off-screen scrolling, and returns to the result page.
+Instead, we focus on query and document content, document position and click label as our training data.
 We measure ranking performance on the annotated set using discounted cumulative gain (DCG) at different truncation levels and mean reciprocal rank (MRR) at $10$, as well as negative log-likelihood (NLL) for click prediction.
@@
 [^10] https://huggingface.co/docs/transformers/model_doc/bert#transformers.FlaxBertForPreTraining The input to our BERT models is: [CLS] query [SEP] doc_title [SEP] doc_abstract.
-Following previous work on Baidu-ULTR (Zou et al., 2022; Chen et al., 2023b), we truncate the input to a maximum length of 128 tokens.
+Following (Zou et al., 2022; Chen et al., 2023b), we truncate the input to a maximum length of 128 tokens.
 Compared to the original BERT training scheme (Devlin et al., 2019), we keep the masked-language modeling task (masking $30\%$ of input tokens at random) but discard the next-sentence prediction task.
@@
 (2023).
-First, we tune our model architecture per dataset based on the pointwise naive model.
+First, we tune our model architecture per set of input features based on the pointwise naive model.
 We tune the number of hidden dimensions $\in\{64,128,256,512,1024\}$ and the number of layers $\in\{2,3,4,5\}$ over three random seeds.
 As both network depth and width can interact with the learning rate, we tune each parameter combination over three learning rates $\in\{0.001,0.0005,0.0001\}$, i.e., we treat the learning rate as a nuisance parameter (Godbole et al., 2023).
-We adopt this incremental tuning procedure as we found no major discrepancies in model architecture between ULTR methods but instead between datasets.
+We adopt this incremental tuning procedure as we found no major discrepancies in model architecture between ULTR methods but instead between sets of input features.
 Subsequently, we adopt the architecture of the pointwise naive model per dataset for all other methods.
@@
 Given the final hyperparameters for each method, we train all methods for 50 epochs, stopping early after five epochs of no improvement of the validation loss computed on clicks.
-Lastly, we highlight that while this tuning step improved the performance of all methods, our main findings are consistent across many different hyperparameter combinations.
+Lastly, while this tuning step improved the performance of all methods, our main findings are consistent across many different hyperparameter combinations.
 6. Results
 Figure 3.
-Comparing ULTR method on pre-trained BERT embeddings and LTR features.
+Comparing ULTR methods on pre-trained BERT embeddings and LTR features.
 We display the average ranking performance measured in DCG@10 over five independent runs and plot a bootstrapped 95% confidence interval.
@@
 Methods based on LambdaRank outperform methods based on the listwise softmax loss, which outperform those based on the pointwise binary cross-entropy loss.
-In contrast, applying ULTR techniques yield marginal improvements at most compared to their naive, non-debiasing counterpart.
+In contrast, applying ULTR techniques yields marginal ranking improvements at most compared to their naive, non-debiasing counterpart.
 In detail, we observe that the pointwise two-tower and IPS models do not consistently and significantly improve performance compared to the naive pointwise loss – IPS is even significantly worse on both BERT features – and RegressionEM is the only pointwise ULTR method that brings small but sometimes significant improvements.
@@
 These results are broadly consistent across input features, suggesting that standard ULTR methods struggle to bring improvement to Baidu-ULTR, regardless of the query-document features.
-Overall, we can answer research questions (RQ1) and (RQ2): on the Baidu-ULTR dataset, unbiased learning-to-rank techniques do not bring any clear benefits, particularly when contrasted with the significant and reliable differences based on the choice of query-document features and ranking loss.
-Our reranking results confirm the initial findings of Zou et al.
+Overall, we can answer research questions (RQ1) and (RQ2): on the Baidu-ULTR dataset, unbiased learning-to-rank methods do not consistently improve ranking performance on expert annotations, particularly when contrasted with the significant and reliable differences based on the choice of query-document features and ranking loss.
+Our reranking results confirm the findings of Zou et al.
 (2022).
@@
 Table 2 gives an overview of applying ULTR methods during language model training.
-In contrast to the inefficacy of ULTR-based rerankers, we observe stark differences due to applying ULTR methods when training language models.
+In contrast to the inefficacy of ULTR-based rerankers, we observe stark differences when applying ULTR methods during language model training.
 Similar to the reranking task, approaches based on the listwise softmax loss perform better than those based on pointwise binary cross-entropy.
@@
 In addition to evaluating ranking performance, we display the negative log-likelihood (NLL) of predicted click scores in Figure 4.
-We restrict this evaluation to pointwise ULTR methods, making well-defined click predictions.
+We restrict this evaluation to pointwise ULTR reranking methods, making well-defined click predictions.
 All ULTR methods robustly improve click prediction compared to a naive loss, with the two-tower model showing the largest improvement on all three datasets.
@@
 These concurrent results suggest that click prediction and ranking performance on annotations are diverging objectives on Baidu-ULTR and training models on clicks does not guarantee improvements in ranking metrics.
-Figure 4. Click prediction performance of pointwise methods measured in negative log-likelihood, lower is better.
+Figure 4. Click prediction performance of pointwise methods measured in negative log-likelihood; lower is better.
 7. Discussion
@@
 Pointwise Naive $7.251$ $0.227$
-Pointwise Two-Tower $7.456$ $0.217$
-Pointwise IPS $6.296$ $0.317$
+Pointwise Two-Tower $7.456$ $0.218$
+Pointwise IPS $6.296$ $0.222$
 Listwise Naive $8.478$
@@
 More complex user behavior.
-Another hypothesis, initially suggested by Zou et al.
-(2022), is that the actual user model is more complicated than the simple PBM we consider in this work.
-This possibility might explain why listwise approaches bring such considerable improvements over pointwise methods and would not be surprising as user studies have identified more complex biases (Agarwal et al., 2019a; Zhuang et al., 2021; Sarvi et al., 2023).
-However, we point out that ULTR methods bringing no benefits would mean that the substantial position bias we identified is negligible against other types of biases.
+Zou et al.
+(2022) suggest that the actual user model might be more complicated than the PBM considered in this work.
+This hypothesis might explain why listwise approaches bring such considerable improvements over pointwise methods and would not be surprising as user studies have identified more complex biases (Agarwal et al., 2019a; Zhuang et al., 2021; Sarvi et al., 2023).
+However, ULTR methods bringing no benefits would mean that the substantial position bias we identified is negligible against other types of biases.
 This assumption, in turn, would contradict related work where PBM-based models trained on more complex user behaviors still improved performance, albeit by not as much as the correct user model (Vardasbi et al., 2020a; Deffayet et al., 2023b).
@@
 Finally, the divergence of click prediction and ranking performance hints at a deeper issue: annotators may not find the same documents relevant as users scrolling through a result page, which might be personalized and part of a multi-query session to find actionable information for the current user needs.
-We hope more user studies or parallel datasets of clicks and expert annotations can clarify this hypothesis.
+We hope further analysis of Baidu-ULTR or parallel datasets of clicks and expert annotations can clarify this hypothesis.
 […]
@@
 7.3. Implications for the field
-Our results starkly contrast with many studies in semi-synthetic simulation (Joachims et al., 2017; Ai et al., 2021; Tran et al., 2021; Saito et al., 2020; Ai et al., 2018a; Hu et al., 2019).
-They call for adjusting these semi-synthetic experimental setups to better reflect real-world challenges.
+Our results starkly contrast with many studies in semi-synthetic simulation (Joachims et al., 2017; Ai et al., 2021; Tran et al., 2021; Saito et al., 2020; Ai et al., 2018a; Hu et al., 2019) and call for adjusting semi-synthetic experimental setups to reflect real-world challenges better.
 The richness of this dataset allows for seeding simulations with more plausible data and, therefore, might bridge the gap between simulations and reality.
@@
 As the largest publicly available dataset comprising both click logs from a major search engine and expert annotations, this dataset constitutes a rare opportunity to assess the progress of click-based learning-to-rank, and especially unbiased learning-to-rank.
-Our main conclusion, however, is that while we have observed substantial improvements relating to the choice of query-document representations and ranking loss (e.g., pointwise or listwise), conventional unbiased learning-to-rank techniques do not bring any clear benefits when applied to the Baidu-ULTR dataset.
+Our main conclusion, however, is that while we have observed substantial improvements relating to the choice of query-document representations and ranking loss (e.g., pointwise or listwise), conventional unbiased learning-to-rank techniques do not bring clear improvements in ranking performance on the Baidu-ULTR dataset.
 Our findings confirm the work by Zou et al.
@@
 Acknowledgements.
-We thank Lixin Zou for his valuable insights and clarifications on the original Baidu work, as well as Jiaxin Mao and Zechun Niu for sharing their learnings from NTCIR 17.
-This research was supported by the Mercury Machine Learning Lab, created by TU Delft, the University of Amsterdam, and Booking.com.
-Maarten de Rijke was supported by the Hybrid Intelligence Center, a 10-year program funded by the Dutch Ministry of Education, Culture and Science through the Netherlands Organisation for Scientific Research, https://hybrid-intelligence-centre.nl.
+We are grateful to Lixin Zou for his valuable insights and clarifications on the original Baidu work, as well as to Jiaxin Mao and Zechun Niu for sharing their learnings from NTCIR 17.
+This research was supported by the Mercury Machine Learning Lab, created by TU Delft, the University of Amsterdam, and funded by Booking.com.
+Maarten de Rijke was supported by the Hybrid Intelligence Center, a 10-year program funded by the Dutch Ministry of Education, Culture and Science through the Dutch Research Council, https://hybrid-intelligence-centre.nl.
 All content represents the opinion of the authors, which is not necessarily shared or endorsed by their respective employers and/or sponsors.